OpenStudy (anonymous):

Solve for x using the quadratic equation. x^2-10x+50=0 The solution set is: { , } Note: If someone can tell me how to simplify the quadratic equation, that would be great since I can't seem to simplify correctly after I plug a, b, and c in.

OpenStudy (anonymous):

Apparently the answer is {5+5i, 5-5i} but I don't know why...

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two rots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

what are your a,b,c ?

OpenStudy (anonymous):

a=1, b=-10, c=50

hartnn (hartnn):

so what is b^2 - 4ac ?

OpenStudy (anonymous):

-300

OpenStudy (anonymous):

this is of the form, ax^2 +bx+c =0 a=1 b=-10 c=50 and the roots are x= +10+/-root{100-4.1.50}/2.1 x= 10+/- root{100-200}/2 10+/- root{-100}/2 10+/- 10i/2 5+/-5i becoz rooot{-1} is treated as i

hartnn (hartnn):

and yes, the solution is 5+/-5i 100-200 = -100 not -300

hartnn (hartnn):

so @Jalberra your b^2-4ac was incorrect, did u get your mistake ? now can u solve on your own ?

OpenStudy (anonymous):

Yes. I got it now. I didn't make b^2 positive which messed everything up. Thanks!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

I was looking at this problem again and b^2 is (-10)^2 which is -100 so how did it become positive?

hartnn (hartnn):

(-10)^2 =(-10)(-10) = +100 minus * minus = plus

OpenStudy (anonymous):

No wonder, I was entering it into my calculator without the parenthesis... Figures, thanks again... so many minor mistakes... -.-

hartnn (hartnn):

happens with all of us sometimes....